The colorless solid phosphorus pentachloride reacts violently with water to give a mixture of phosphoric acid and hydrochloric acid .

(a) Write a balanced chemical equation for this reaction.

_______+________--_______+_______

(b) Determine the concentration (in moles per liter) of each of the acids that result from the complete reaction of 3.57×10-2 moles of phosphorus pentachloride with enough water to give a solution volume of 806 mL.

phosphoric acid : ______ M

hydrochloric acid: _______ M
1
Expert's answer
2019-03-13T00:49:43-0400

a) PCl5 + 4H2O ->H3PO4+5HCl

b) According to equation mole ratio n(PCl5):n(H3PO4)= 1:1, then n(H3PO4)=n(PCl5)=3.57*10-2mol

c(H3PO4)=3.57*10-2mol/0.806L=0.0443mol/L= 0.0443M

According to equation mole ratio n(PCl5):n(HCl)=1:5, then n(HCl)=5*n(PCl5)=5*3.57*10-2=0.179mol

c(HCl)=0.179mol/0.806L=0.221mol/L=0.221M.